OpenStudy (anonymous):

Can you tell me the steps I took to solve the problem below. 2ax + aby + a / a 2ax + aby + a a a a 2 · ax + aby + a 1 a a a 2x+by+a

OpenStudy (anonymous):

thats right :)

OpenStudy (anonymous):

but what are the steps?

OpenStudy (anonymous):

what is the question?? (2ax + aby + a )/ a ????

OpenStudy (anonymous):

oh i thought you wanted your steps checked >.< my bad.

OpenStudy (chaise):

Are we solving for x?

OpenStudy (anonymous):

no it is polynoimal, which the teacher did the steps to solve but wants us to fiqure out

OpenStudy (anonymous):

The steps are pretty much there. You start with: \[\frac{2ax+aby+a}{a}\]Then you split the division up: \[\frac{2ax}{a}+\frac{aby}{a}+\frac{a}{a}\]Using the fact that a/a = 1, we now have: \[2x*\frac{a}{a}+by*\frac{a}{a}+\frac{a}{a} = 2x+by+1\]

OpenStudy (anonymous):

thank you for your help.